A 68.4 g drink is left on a turntable causing it to turn with an angular speed of 8.80 rad/s. If the drink has coefficients of friction of μs = 0.960 and μk = 0.450 with the turntable, what is the furthest distance that the drink could have been placed from the centre of the turntable for the drink not to slide?

a) 0.121 m

b) 0.00133 m

c) 0.0569 m

d) 0.0124 m

___________
Answer #1

Here, Radial Force should be equal to frictional force.

mw^2r = umg

8.8^2*r = 0.960*9.8

   r= 0.960*9.8/8.8^2

    = 0.121 m

Option a is correct